OpenStudy (anonymous):

The ratio of the weight of an earth to the weight of the same object on Pluto is 100 to 3 if a buffalo weight 3858 pounds on earth find the buffalo's weight on Pluto?

Directrix (directrix):

I'm thinking that the following proportion will give the buffalo's weight on Pluto: 100/3 = 3858/x where x is the animal's weight on Pluto 100*x = 3858 * 3 x = ? @Gafra

OpenStudy (anonymous):

I did it this way and I got the answer thanks

Directrix (directrix):

You are welcome.